OpenStudy (anonymous):

Find the equation of the plane perpendicular to the curve x=t^5, y=2t^2, z=3t at t=1

OpenStudy (anonymous):

First find the point at t=1 (1,2,2)

OpenStudy (anonymous):

then find the tangent vector \[ (x'(t), y'(t), z'(t))= ( 5 t^4, 4 t, 3) \\ at \quad t=1 \\ (5,4,3) \]

OpenStudy (anonymous):

The equation of the plane is\[ 5(x-1) + 4(y-2) + 3(z-2)=0 \]

OpenStudy (anonymous):

THANK YOU!